suppose you're going to make a password that consists of 5 characters chosen from {1, 2, 4, 9, d, i, k, m, n, w, z} How many different passwords can you make if you cannot use any character more than once in each password?
Answer #1

number of different passwords =N(11 choices for first, 10 for 2nd,,,,,,,,,) =11*10*9*8*7=55440